Write the standard form of the equation of the line with slope -4 passing through the point (2,-5). Can some help
Answer by Prithwis(166)

Slope intercept form is y = mx + c where m is the slope of the line.
So, y = -4x+c
Passes thru (2,-5): -5 = -8+c => c = 3;
Equation is y = -4x+3
When converted to standard form: 4x+y = 3